OpenStudy (charlotte123):

Triangle EFG is similar to Triangle QRS. Gary builds the proportion below to solve for the missing side QR. EF/FG = QS/QR Explain what mistake Gary made and how you would correct it.

Directrix (directrix):

Triangle EFG is similar to Triangle QRS means that corresponding vertices are E and Q, F and R, and G and S. Lengths of corresponding sides of similar triangles are in proportion so: EF/QR = FG/RS = EG /QS ------- Gary wrote EF/FG = QS/QR. That is equivalent to the proportion: EF/QS = FG /QR Those are not corresponding sides of these two similar triangles. Gary should have followed the correspondences given by the statement of the similar triangles and written EF/QR = FG/RS.

jhonyy9 (jhonyy9):

EF*QR =QS*FG QR=(QS*FG)/EF

OpenStudy (charlotte123):

Thanks Directrix but are you sure that is the right answer?:P